OpenStudy (anonymous):

What value completes the square for the expression? show work? x2 + 14x

OpenStudy (anonymous):

x^2 + bx -> take half of b and then square that. (b/2)^2

OpenStudy (anonymous):

In this example, b = 14

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

Take 1/2 of the number in front of the x (14) and square it. So 1/2 of 14 is 7 and 7 squared is 49 x^2 + 14x + 49 = (x + 7)(x + 7) = (x + 7)^2 Answer: 49